OpenStudy (anonymous):

Write a proportion for the following. When surveyed, five out of every eight students liked the food in the cafeteria. If 60 liked the food, how many were surveyed? 5/8 = x/60 5/8 = 60/x 5/60 = x/8

OpenStudy (anonymous):

5/8=x/60?

OpenStudy (apoorvk):

Uh-oh. The denominator represents the no. of people who were surveyed, so shouldn't 'x' be the denominator and '60' the numerator? What do you say?

OpenStudy (anonymous):

how you do that?

OpenStudy (apoorvk):

Umm, 5 *out of* 8 liked, so 60 *out of* 'x' would like it right? So, that would be: \[ \frac58 = \frac {60}{x}\] Understanding this?

OpenStudy (anonymous):

wah da hell -.- wonder i got this thank you ! 96 right?

OpenStudy (apoorvk):

Da Hell '96'!! Awwyeah you're right!!!!!!